# Sequential Kernelized Stein Discrepancy

### v1

@@ A.1 Comparison to batch setting
Figure 2: Proportion of rejections for the Gaussian distribution considered in Section 6 under three different alternatives with (I) $\theta=0.5$, (II) $\theta=0.75$, (III) $\theta=1$. We emphasize that the (proposed) sequential test always ends up rejecting the null hypotheses, while the batch test will not do so if the original sample size is too small.
A.2 Empirical verification of the lower bound of the wealth
@@
Figure 3: Average empirical stopping values $\tau$, alongside their $95\%$ empirical confidence intervals, as a function of $r^{*}$ (defined in Theorem 2) in the Gaussian setting, with $\alpha=0.05$, across $500$ simulations, for the LBOW and aGRAPA strategies. We emphasize that all the stopping times are dominated by the approximate upper bound $log(1/\alpha)/r^{*}$.
Appendix B Proofs
B.1 Notation
@@ B.3 Auxiliary results
$\leq\sup_{i\geq N_{1}}\left\{\left\|\frac{1}{i-1}\sum_{k=1}^{i-1}\xi_{p}(\cdot,X_{k})-\mathbb{E}[\xi_{p}(\cdot,X)]\right\|_{\mathcal{H}^{d}}\right\}\sup_{t\geq 2N_{1}}\left\{\frac{1}{t-N_{1}}\sum_{i=N_{1}}^{t}\left\|\xi_{p}(\cdot,X_{i})\right\|_{\mathcal{H}^{d}}\right\}$
@@
B.4 Proofs of main theorems
@@
and so $\limsup_{t\to\infty}\frac{1}{t}\sum_{i=1}^{t}\left|g_{i}(X_{i})\right|\leq\mathbb{E}[\left|g^{*}(X)\right|]$. Furthermore, we now that $\frac{1}{t}\sum_{i=1}^{t}g_{i}^{2}(X_{i})\stackrel{{\scriptstyle a.s.}}{{\to}}\mathbb{E}[\left(g^{*}(X)\right)^{2}]$. Given that $\lambda\stackrel{{\scriptstyle a.s.}}{{\to}}\lambda^{*}$ and $\kappa$ is continuous, it follows that $\kappa(\lambda)\stackrel{{\scriptstyle a.s.}}{{\to}}\kappa(\lambda^{*})$. Thus, by the SLLN, there exist $B>0$ and $N_{1}\in\mathbb{N}$ such that

### v2

@@ A.3 Comparison to batch setting
Figure 5: Proportion of rejections for $500$ simulations for the Gaussian distribution considered in Section 6 under three different alternatives with (I) $\theta=0.5$, (II) $\theta=0.75$, (III) $\theta=1$. We emphasize that the (proposed) sequential test always ends up rejecting the null hypotheses, while the batch test will not do so if the original sample size is too small.
We would like to emphasize the lack of anytime validity of the classical, batch setting test. Figure 6 displays the proportion of rejections for both the proposed sequential test, and the batch test run sequentially (once every time a new observation is observed). Not surprisingly, the batch test rejection rates go well above the desired $\alpha=0.05$ type-I error. This example illustrates the lack of anytime validity of classical test.
Figure 6: Proportion of rejections for $100$ simulations for the Gaussian distribution considered in Section 6 under the null for the proposed sequential test, and the batch test run sequentially (once every time a new observation is observed). While the sequential test preserves anytime valid guarantees, the ‘sequentialized’ batch test lacks type-I error control.
A.4 Composite null hypotheses
Section 4.2 presented the extension of the proposed test to composite null hypotheses. We devote this appendix to illustrating the validity and power of such an extension. Let the null hypothesis be $H_{0}:Q\in\mathcal{P}$ against $H_{1}:Q\not\in\mathcal{P}$. Throughout, the null hypotheses will be combinations of the Gaussian-Bernoulli restricted Boltzmann machines (RBMs) from Section 5.
In particular (and analogously to Appendix A.1), we take $d_{h}=10$, and $d=50$. We sample from such RBMs using Gibbs sampling with a burn-in of 1000 iterations. Throughout, $Q$ is taken as an RBM with $b=0$, $c=0$, and matrix $B$ such that $B_{ij}$ is one if visible node $i$ is connected to hidden node $j$, and zero otherwise, with each hidden node connected to five visible nodes. For two different experiments, we consider the two null hypotheses $\mathcal{P}=\{Q,Q_{a,1}\}$ and $\mathcal{P}=\{Q_{a,1},Q_{a,2}\}$, where $Q_{a,1}$ is an RBM with each entry of $B$ is shifted by $0.5$ and $Q_{a,2}$ is taken with $b=1$. Figure 7 displays the logarithmic wealth processes for these two scenarios. Note that the logarithmic wealth process never crosses the threshold $\log(1/0.05)$ in the first case (showing the empirical anytime validity of the test), while the latter case shows comparable power to those experiments presented in Section A.1.
Figure 7: Average logarithmic wealth for $1000$ simulations for the composite null hypothesis (I) $\mathcal{P}=\{Q,Q_{a,1}\}$, (II) $\mathcal{P}=\{Q_{a,1},Q_{a,2}\}$, where $Q$ is the RBM defined in Section 5 with $b=0$, $c=0$, and matrix $B$ such that $B_{ij}$ is one if visible node $i$ is connected to hidden node $j$, and zero otherwise, with each hidden node connected to five visible nodes; $Q_{a,1}$ is the same RBM with each entry of $B$ is shifted by $0.5$, and $Q_{a,2}$ is taken with $b=1$.
A.5 On the tightness of the bounds and statistical power
In Section 5, we carefully exhibited how to derive sensible bounds $M_{p}$ for different families of distributions. Here, we explore how the tightness of the bounds affects the statistical power of the test. Figure 8 displays the wealth processes for the Gaussian distribution from Section 5 under three different alternatives, using the original bound obtained in Section 5 alongside looser bounds. We highlight that all logarithmic wealth processes are linear (i.e., the wealth growth is exponential), but looser bounds lead to smaller slopes.
Figure 8: Average logarithmic wealth for $1000$ simulations for the Gaussian distribution when using bounds of the form $k\cdot M(X)$, where $M(X)$ is the original bound derived in Section 5 and $k\in\{1,2,3,4\}$, under the alternatives (I) $\theta_{1}=0.5$, (II) $\theta_{1}=0.75$, (III) $\theta_{1}=1$. It can be clearly seen that all the logarithmic wealth processes are linear (i.e., the wealth growth is exponential), but higher multiplicative factors $k$ (i.e., looser bounds) have smaller slopes.
A.6 Empirical Verification of the Lower Bound of the Wealth
@@
Figure 9: Average empirical stopping values $\tau$, alongside their $95\%$ empirical confidence intervals for $1000$ simulations, as a function of $r^{*}$ (defined in Theorem 2) in the Gaussian setting, with $\alpha=0.05$, across $500$ simulations, for the LBOW and aGRAPA strategies. We emphasize that all the stopping times are dominated by the approximate upper bound $log(1/\alpha)/r^{*}$.
Appendix B Comparison to related work
We present in this appendix a comprehensive comparison of our contribution and that of Zhou and Liu, (2024), which also proposed a sequential KSD goodness-of-fit test that is based on a supermartingale construction and Ville’s inequality.
The primary distinction between the contributions lies in the uniform boundedness assumption, which is adopted in Zhou and Liu, (2024) but deliberately avoided in our work. Specifically, their work assumed that the scores are uniformly bounded by 1 (Zhou and Liu,, 2024, Assumption 2.4). We would like to emphasize that this is, in principle, a rather restrictive assumption. For instance, even the simple case of a Gaussian distribution with an IMQ kernel does not satisfy this assumption. Thus, our method is strictly more general and can accommodate more scenarios, e.g., the three experiments presented in Section 6 could not have been addressed using their procedure.
In principle, there could be other alternatives to work around unbounded score functions. For instance, one could work with tilted base kernels that make the Stein kernel bounded even if the score function is not bounded (Barp et al.,, 2024, Theorem 7). Nonetheless, current results such as Barp et al., (2024, Theorem 7) still require bounds to ‘tilt’ the base kernel, but do not provide such bounds (the authors’ goal is solely to prove that boundedness of the Stein kernel may be assumed without loss of generality in their setting, without further interest in the bound itself). One of the main contributions of our paper is to show how to obtain tight workable bounds, and illustrate that such derivations yield powerful tests. In order to potentially use Barp et al., (2024, Theorem 7) to obtain sequential tests, one would have to carefully study the actual bounds derived from such a theorem. This would eventually lead to an analysis very similar to the one presented throughout this contribution.
The lack of uniform boundedness also conveys a number of deep theoretical challenges. The theoretical foundation of Zhou and Liu, (2024) is analogous to the one presented in Podkopaev et al., (2023). See e.g. Zhou and Liu, (2024, Theorem 2.5) and Podkopaev et al., (2023, Theorem 2.4), alongside their proofs. However, the lack of uniform boundedness makes such proofs break in (at least) two different places: (i) after applying the Cauchy-Schwarz inequality, see Zhou and Liu, (2024, Equation (47)), (ii) the analysis of the Online Newton Step (ONS) strategy for selecting betting fractions. In order to circumvent the above challenges, this contribution replaces the ONS strategy for aGRAPA/LBOW strategies and presents novel theoretical guarantees, which are exhibited in Appendix C.3.555Some of the arguments presented in Appendix C.3 may be of independent interest. To the best of our knowledge, no other contribution in the ‘testing by betting’ literature exploits outcomes that are lower bounded but not upper bounded. Lastly, we have experimentally found that the aGRAPA and LBOW strategies proposed in our contribution yield substantially more powerful tests than the ONS strategy proposed in Zhou and Liu, (2024) in all the cases considered. As exhibited in Figure 10, the average log wealth at round $100$ for aGRAPA roughly doubles the log wealth achieved by the ONS strategy for all the experiments.
Figure 10: Average logarithmic wealth alongside $95\%$ empirical confidence intervals for $1000$ simulations under the alternative for (I) the Gaussian distribution, (II) the intractable model, (III) the restricted Boltzmann machine with shifted $B$, (IV) the restricted Boltzmann machine with bias $b=1$. We highlight that the logarithmic wealth achieved by ONS is roughly doubled by that obtained by the aGRAPA strategy.
Appendix C PROOFS
C.1 Notation
@@ C.3 Auxiliary Results
$\leq\sup_{i\geq N_{1}}\left\{\left\|\frac{1}{i-1}\sum_{k=1}^{i-1}\xi_{p}(\cdot,X_{k})-\mathbb{E}[\xi_{p}(\cdot,X)]\right\|_{\mathcal{H}^{d}}\right\}$
×$\quad\times\sup_{t\geq 2N_{1}}\left\{\frac{1}{t-N_{1}}\sum_{i=N_{1}}^{t}\left\|\xi_{p}(\cdot,X_{i})\right\|_{\mathcal{H}^{d}}\right\}$
@@
C.4 Proofs of Main Theorems
@@
and so $\limsup_{t\to\infty}\frac{1}{t}\sum_{i=1}^{t}\left|g_{i}(X_{i})\right|\leq\mathbb{E}[\left|g^{*}(X)\right|]$. Furthermore, we know that $\frac{1}{t}\sum_{i=1}^{t}g_{i}^{2}(X_{i})\stackrel{{\scriptstyle a.s.}}{{\to}}\mathbb{E}[\left(g^{*}(X)\right)^{2}]$. Given that $\lambda\stackrel{{\scriptstyle a.s.}}{{\to}}\lambda^{*}$ and $\kappa$ is continuous, it follows that $\kappa(\lambda)\stackrel{{\scriptstyle a.s.}}{{\to}}\kappa(\lambda^{*})$. Thus, by the SLLN, there exist $B>0$ and $N_{1}\in\mathbb{N}$ such that